# Task-Oriented Feature Compression for Multimodal Understanding via Device-Edge Co-Inference

### v1

@@ IV-C Selective Entropy Model
where the length of $\boldsymbol{s}_{n,i}$ is equal to the number of entropy models $n_{\rm e}$. During training, all entropy models encode and decode merged features separately, and the final quantized features are the weighted sum of the decoding results from each entropy model, given by:
@@
$R=-\frac{1}{n_{\rm p}n_{\rm c}d_{\rm v}}\sum_{e=1}^{n_{\rm e}}w_{n,i}^{(e)}(\log_{2}p(\bar{\boldsymbol{y}}^{(e)}|\bar{\boldsymbol{z}}^{(e)}+\log_{2}p(\bar{\boldsymbol{z}}^{(e)})).$ (14)
@@
where $T$ is the temperature that controls the distribution of the weighting coefficients. At the beginning of training, the temperature is high, and the weighting coefficients of different entropy models are similar, allowing all entropy models to learn the basic coding capabilities. As training progresses, the temperature gradually decreases, and the weighting coefficients converge towards the single optimal selection employed during inference, which enables the specialization of different entropy models. To avoid the scenario where certain entropy models are never selected, we introduce balance loss to the RD loss in (9), expressed as:
@@
where $\alpha$ is the weight of the balance loss. A high value of $\alpha$ encourages a uniform usage frequency across experts in each batch but hinders the specialization of each entropy model. Thus, the value of $\alpha$ is carefully selected to maximize the compression efficiency. The detailed training procedures are summarized in Algorithm 1.
@@
During inference, each compressed feature is encoded and decoded using the entropy model with the highest score, which is equivalent to temperature approaching zero. The weighting coefficient is expressed as:
@@ V-A1 Network Implementation
We employ LLaVA-OneVision-7B [7] as the backbone model and SigLIP-SO400M [10] as the vision encoder. The resolution of each image patch is $384\times 384$ pixels, the number of visual features is $n_{\rm v}=729$, and the dimension of each visual feature is $d_{\rm v}=1152$. The number of merged features $n_{\rm c}$ varies from 8 to 16, and the number of entropy models is $n_{\rm e}=16$. The learnable entropy models are implemented based on the CompressAI package [50]. The structures of each learnable entropy model and the router network are illustrated in Figs. 3 and 5, respectively.
@@ V-A2 Training Details
The training data consists of the instruction tuning data utilized in the second training stage of LLaVA-1.5 [51], accounding for a small portion of data used in backbone model training. The proposed model is trained for only one epoch, with the trainer settings identical to those used in the second training stage of the backbone model. With the rank of LoRA adapters [49] set to 64, the total number of trainable parameters is approximately 300 million. The Lagrange multiplier $\lambda$ varies between 0.003 and 0.03 to achieve different trade-offs between data transmission overhead and the performance of downstream tasks The weight of the balance loss is set as $\alpha=0.001$. Additionally, the temperature of softmax function $T$ exponentially decreases from 10 to 0.1 throughout the training process.
@@ V-A3 Evaluation Details
The image understanding capability is measured by seven VQA benchmarks, including RealWorldQA, the perception score of MME [52], the test set of AI2D [53], the English development set of MMBench [54], MMStar [55], the test set of ScienceQA [56], and the validation set of MMMU [57]. These benchmarks sufficiently evaluate multiple aspects of multimodal understanding, including but not limited to reasoning, perception, mathematics, translation, and common sense. The implementation of evaluation is based on the VLMEvalKit package [58]. The user device utilized for testing is an NVIDIA Jetson AGX Orin, and the edge server is equipped with an adequate number of NVIDIA RTX 4090 GPUs. We compare the proposed method against two data-oriented image compression methods: JPEG [20] and WebP [21]. To ensure a fair comparison in terms of system latency, we train a model with the proposed FMM using the distortion loss in (10). This model is utilized when measuring the rate-performance curves of traditional image compression methods: JPEG and WebP.
@@ V-B1 Rate-Performance Curve
We train multiple instances of proposed TOFC models with varying Lagrange multipliers $\lambda$ and numbers of merged features $n_{\rm c}$ to explore the trade-off between data transmission overhead and task performance. For traditional JPEG and WebP compression methods, we adjust their quality settings to obtain the rate-performance curves. Additionally, we integrate these traditional compression methods with the FMM proposed in Section IV-A to accelerate LLM inference at the edge server, thereby ensuring similar system latency compared with the proposed TOFC method. Fig. 6 presents the rate-performance curves of different methods on RealWorldQA and MME benchmarks. For the RealWorldQA benchmark with high-resolution images, the proposed TOFC method achieves 25% to 45% reductions in data transmission across various target task performances in comparison to WebP compression without FMM. The reduction is even more pronounced when compared to JPEG compression without FMM, ranging from 35% to 60%. When FMM is incorporated to match the system latency of the TOFC method, traditional WebP and JPEG compressions struggle to maintain comparable task performance, incurring 5.9$\times$ and 6.7$\times$ data transmission overhead, respectively. For the MME benchmark with low-resolution images, the proposed TOFC method still provides approximately 25% and 42% reductions in data transmission, compared to WebP and JPEG compressions without FMM, respectively. When FMM is applied to accelerate inference, traditional WebP and JPEG compressions require 3.2$\times$ and 3.6$\times$ data transmission, respectively, to achieve the same task performance as the TOFC method. These results demonstrate the effectiveness of the TOFC method in discarding redundant information in visual features while maintaining task performance.
Fig. 7 demonstrates the rate-performance curves of different methods on the seven VQA benchmarks elaborated in Subsection V-A-3. The task performance is assessed based on the average performance across all benchmarks, normalized by the performance of the 7B backbone. The proposed TOFC method with $n_{\rm c}=8$ and $n_{\rm c}=12$ achieves approximately 55% and 45% reductions in data transmission, respectively, compared with the combination of WebP and FMM under identical $n_{\rm c}$ settings. In comparison to JPEG coupled with FMM, the TOFC method with $n_{\rm c}=8$ and $n_{\rm c}=12$ provides even more substantial data size savings of 68% and 54%, respectively. These results underscore the efficiency of the proposed TOFC method across a variety of VQA benchmarks. Our TOFC method minimizes data transmission overhead while maintaining high task performance, thereby enhancing user experience and enabling timely response for delay-sensitive tasks.
@@ V-B2 Latency Analysis
Figure 8: Average inference latency per user request under different communication rates. The total latency consists of on-device computation time, data transmission latency, and server-side computation time. (a) and (b) show results on RealWorldQA and MME benchmarks, respectively.
@@
To evaluate the low latency advantage of our TOFC, we conduct a comparison of the end-to-end system latency across different methods, where the total latency consists of on-device computation time, data transmission latency, and server-side computation time. To ensure a fair comparison, the Lagrange multiplier $\lambda$ for the proposed model and the quality settings for traditional methods are carefully selected to achieve similar task performance. Fig. 8 demonstrates the average inference latency per user request under different communication rates on the RealWorldQA and MME benchmarks. For the RealWorldQA benchmark with high-resolution images, the proposed TOFC method achieves approximately a 50% reduction in system latency compared with traditional image compression methods, JPEG and WebP, under low communication rates. At higher communication rates, the latency reduction becomes more prominent, as the TOFC method significantly accelerates LLM inference at the edge server by reducing the number of visual tokens. The system latency of WebP compression is slightly higher than that of JPEG compression, as the increased on-device computation time outweighs the reduced data transmission latency. These results demonstrate the superior latency performance of the proposed TOFC method compared to data-oriented image compressions, benefiting from the proposed FMM method, which significantly reduces computation time at the edge server. Without the reduction in data transmission, traditional server-only and device-only inference schemes both suffer from excessive latencies of over 3 seconds, causing severe degradations on the user experience. Moreover, even at high communication rates, the data transmission latency is 1.8 seconds on average, constituting the majority of system latency in mainstream server-only inference schemes. These results highlight the necessity of minimizing data transmission overhead especially when the resolution of the input image is high. For the MME benchmark with low-resolution images, the proposed TOFC method manages to complete inference in less than half the time compared to JPEG and WebP compressions. The system latency of server-only inference ranges between 0.64 seconds and 1.53 seconds, lower than that observed in the RealWorldQA benchmark due to the lower image resolution. However, the TOFC method with $n_{\rm c}=16$ reduces inference time to only one-third of that required by server-only inference, without noticeable degradation in task performance.
@@ V-C1 Feature Merging Methods
Methods AI2D MMBench MME MMStar RealWorldQA Avg.
Proposed FMM 74.61 78.78 1540.8 50.33 63.92 91.55%
FMM w/o merging 73.83 77.06 1484.4 48.27 61.18 88.77%
VisionZip [27] 70.85 74.23 1484.8 46.86 51.37 84.05%
VisionZip merging [27] 71.99 75.34 1525.9 49.55 57.25 87.70%
Q-Former [59] 68.69 68.04 1379.7 42.83 53.99 80.15%
@@ V-C2 Entropy Coding Methods
To verify the effectiveness of the proposed learnable and selective entropy model with hyperprior in reducing data transmission overhead while maintaining task performance, we conduct an ablation study on alternative entropy coding methods. Specifically, we replace the entropy coding method in the proposed TOFC method with the following three methods. First, we discard the selective entropy model architecture proposed in Section IV-C, retaining only a single entropy model. Second, we substitute the learnable entropy model with a heuristic Laplacian probability model, whose mean and scale are calculated from the statistics of the quantized features in each batch. A linear scaling is applied before quantization, such that the 99.9% percentile of the merged features is mapped to the maximum integer. Third, the learnable factorized density model, originally used for estimating the probability distribution of the hyperprior $\boldsymbol{z}$, is directly employed to model the merged features $\boldsymbol{y}$. For the first and third learnable methods, we train multiple models with varying Lagrange multipliers $\lambda$ and numbers of merged features $n_{\rm c}$ to obtain the rate-performance curve. For the second heuristic method, we adjust the bit length used for quantization and the number of merged features $n_{\rm c}$ to traverse the rate-performance curve.
As shown in Fig. 10, the proposed selective entropy model outperforms alternative entropy coding methods across all data sizes. Compared with a single entropy model, the proposed selective entropy model achieves a 13% to 22% reduction in data transmission while maintaining similar performance. Compared with the heuristic probability model, the proposed entropy model reduces data transmission overhead by approximately 40%, showing that the learnable entropy model provides a more accurate estimation of the statistics of each merged feature. The learnable entropy model without hyperprior delivers poor performance, confirming that the factorized density model alone is insufficient for estimating the probability distribution without the knowledge extracted by the hyperprior.

### v3

@@ IV-C Selective Entropy Model
where the length of $\boldsymbol{s}_{n,i}$ is equal to the number of entropy models $n_{\rm e}$. The idea of adaptive routing and activation of model weights is inspired by the mixture-of-experts (MoE) architecture widely adopted in LLMs [54, 55], which remains unexplored for learnable entropy models. Based on the unique characteristics of entropy models, we employ different strategies in the training and inference processes. During training, all entropy models encode and decode merged features separately, and the final quantized features are the weighted sum of the decoding results from each entropy model, given by:
@@
$R=-\frac{1}{n_{\rm p}n_{\rm c}d_{\rm v}}\sum_{e=1}^{n_{\rm e}}w_{n,i}^{(e)}(\log p(\bar{\boldsymbol{y}}^{(e)}|\bar{\boldsymbol{z}}^{(e)})+\log p(\bar{\boldsymbol{z}}^{(e)})).$ (16)
@@
where $T$ is the temperature that controls the distribution of the weighting coefficients. The temperature is high at the beginning of training, and the weighting coefficients of different entropy models are similar, allowing all entropy models to learn the basic coding capabilities. The temperature gradually decreases as training progresses, and the weighting coefficients converge towards the single optimal selection employed during inference, which enables the specialization of different entropy models. To avoid the scenario where certain entropy models are never selected, we introduce balance loss to the RD loss in (9), expressed as:
@@
where $\alpha$ is the weight of the balance loss. A high value of $\alpha$ encourages a uniform usage frequency across entropy models in each batch but hinders the specialization of each entropy model. Thus, the value of $\alpha$ is carefully selected via experiments to maximize the compression efficiency. The detailed training procedures are summarized in Algorithm 1.
@@
During inference, each merged feature is encoded and decoded using the entropy model with the highest score to minimize data transmission, which is equivalent to temperature approaching zero. The weighting coefficient is expressed as:
@@ V-A1 Network Implementation
We employ LLaVA-OneVision-7B [7] as the backbone model and SigLIP-SO400M [10] as the vision encoder. The resolution of each image patch is $p\times p=384\times 384$ pixels, and the number of patches in the height and width dimensions are given by $n_{\rm h}=\lceil\frac{h}{p}\rceil$ and $n_{\rm w}=\lceil\frac{w}{p}\rceil$, respectively, where $\lceil\cdot\rceil$ denotes rounding up to the nearest integer. The number of visual features for each patch is $n_{\rm v}=729$, and the dimension of each visual feature is $d_{\rm v}=1152$. The number of merged features $n_{\rm c}$ varies from 8 to 32, and the number of entropy models is $n_{\rm e}=16$. The number of nearest neighbors is fixed at $K=5$, as we find negligible variations in task performance after adjusting its value. The learnable entropy models are implemented based on the CompressAI package [56]. The structures of each learnable entropy model and the router network are illustrated in Figs. 3 and 5, respectively.
@@ V-A2 Training Details
The training data consists of the 665K instruction tuning samples utilized in the second training stage of LLaVA-1.5 [33], accounting for a small portion of 4.8M samples used in backbone model training. The proposed model is trained for only one epoch, with the trainer settings identical to those used in the second training stage of the backbone model [7]. With the rank of LoRA adapters [53] set to 64, the total number of trainable parameters is approximately 300 million, only a small percentage of the total parameter count of 8 billion. The Lagrange multiplier $\lambda$ varies between 0.003 and 0.03 to achieve different trade-offs between data transmission overhead and the performance of downstream tasks. The weight of the balance loss is set as $\alpha=0.001$. Additionally, the temperature of softmax function $T$ exponentially decreases from 10 to 0.1 throughout the training process. The training process takes 8.1 hours on a server with eight RTX 4090 GPUs, substantially lower than the backbone model [7] which is trained for 66 hours in total on 256 A100 GPUs or 128 H100 GPUs.
@@ V-A3 Evaluation Details
The image understanding capability is measured by the following seven VQA benchmarks.
  • The RealWorldQA benchmark evaluates real-world spatial understanding via 765 multiple-choice questions on high-resolution images, which require scene detail recognition and commonsense reasoning.
  • The MME benchmark [57] employs custom-designed, manually annotated instruction-answer pairs to measure both perception and cognition abilities of LMMs across 14 distinct subtasks.
  • The test set of AI2 diagrams (AI2D) [58] is comprised of 3088 grade-school science diagrams with question-answer pairs to evaluate the diagram understanding capabilities.
  • The English development set of MMBench [59] comprises 4329 multiple-choice questions that assess diverse multimodal understanding capabilities of LMMs.
  • The MMStar benchmark [60] collects 1500 human-selected and validated samples, evaluating 6 core capabilities and 18 detailed axes of LMMs.
  • The test set of ScienceQA [61] includes 2017 multiple choice questions in natural science, social science, and linguistics, sourced from elementary and high school science curricula.
  • The validation set of massive multi-discipline multimodal understanding (MMMU) benchmark [62] contains 1050 multimodal questions from college exams, quizzes, and textbooks spanning 30 subjects to evaluate expert-level reasoning capabilities.
The implementation of evaluation is based on the VLMEvalKit package [63]. The user device utilized for testing is an NVIDIA Jetson AGX Orin, and the edge server is equipped with an adequate number of NVIDIA RTX 4090 GPUs. We compare the proposed method against four data-oriented image compression methods, including traditional methods JPEG [21] and WebP [22], and neural codecs hyperprior [24] and ELIC [26]. To ensure a fair comparison in terms of system latency, we train a model with the proposed FMM using the distortion loss in (10). This model is utilized when measuring the rate-performance curves of the four data-oriented image compression methods.
@@ V-B1 Rate-Performance Curve
[…]
Fig. 7 demonstrates the rate-performance curves of different methods on the seven VQA benchmarks elaborated in Subsection V-A-3. The task performance is assessed based on the average performance across all benchmarks, normalized by the performance of the 7B backbone. The proposed TOFC method achieves approximately 23% and 32% reductions in data transmission, compared with ELIC and hyperprior codecs, respectively. In comparison to traditional JPEG and WebP compressions, TOFC provides even more substantial data size savings of 47% and 37%, respectively. These results underscore the efficiency of the proposed TOFC method across a variety of VQA benchmarks. Our TOFC method minimizes data transmission overhead while maintaining high task performance, thereby enhancing user experience and enabling timely response for delay-sensitive tasks.
@@ V-B2 Latency Analysis
Figure 8: Average inference latency per user request (the bar plot corresponding to the left y-axis) and benchmark score (the line plot corresponding to the right y-axis) under different communication rates. The total latency consists of on-device computation time, data transmission latency, and server-side computation time. Subfigures (a) and (b) show results on RealWorldQA and MME benchmarks, respectively.
To evaluate the low latency advantage of our TOFC, we conduct a comparison of the end-to-end system latency across different methods, where the total latency consists of on-device computation time, data transmission latency, and server-side computation time. To ensure a fair comparison, the Lagrange multiplier $\lambda$ for neural codecs and the quality settings for traditional methods are carefully selected to achieve similar task performance. Fig. 8 demonstrates the average inference latency per user request and the benchmark score under different communication rates on the RealWorldQA and MME benchmarks.
@@
For the RealWorldQA benchmark with high-resolution images, the proposed TOFC method achieves system latency reductions between 32% and 47% compared with traditional image compression methods, JPEG and WebP, under various communication rates. The system latency of WebP compression is slightly higher than that of JPEG compression, as the increased on-device computation time outweighs the reduced data transmission latency. Compared with ELIC, TOFC achieves even more significant latency reductions from 43% to 63%, since the complex context calculation in modern neural compression methods incurs excessive on-device computation time. At higher communication rates, the latency reduction becomes more prominent, as TOFC significantly accelerates LLM inference at the edge server by reducing the number of visual tokens. These results demonstrate the superior latency performance of TOFC compared to data-oriented image compressions, benefiting from the proposed FMM method, which significantly reduces computation time at the edge server. As for the traditional single-node computation, both server-only and device-only inference schemes suffer from excessive latencies of over 3 seconds, causing severe degradations on the user experience. Moreover, even at high communication rates, the average data transmission latency is 1.8 seconds without compression, constituting the majority of system latency in mainstream server-only inference schemes. These results highlight the necessity of minimizing data transmission overhead, especially when the resolution of the input image is high.
For the MME benchmark with low-resolution images, the proposed TOFC method manages to complete inference in less than half the time compared to ELIC. Compared with JPEG and WebP compressions, the inference latency of TOFC are approximately 67% and 59% those of the JPEG and WebP compressions. The system latency of server-only inference ranges between 0.64 seconds and 1.53 seconds, lower than that observed in the RealWorldQA benchmark due to the lower image resolution. However, the TOFC method still reduces inference time to only one-third of that required by server-only inference under low communication rate, without noticeable degradation in task performance. As for device-edge co-inference where the uncompressed visual features are directly transmitted to the edge server, the benchmark score is the same as server-only inference with 7B model, but the latency is 40 to 60 times higher, thus not illustrated in Fig. 8.
@@ V-B3 Case Study
V-B4 Extension to other LMMs
To demonstrate that the proposed method can be generalized to other LMMs with alternative vision encoders, we extend our experiments to the LLaVA-1.5-7B [33] model with a CLIP vision encoder [9]. This LMM resizes and pads each input image to a resolution of $p\times p=336\times 336$ as the only input patch ($n_{\rm p}=1$). The number of visual features is $n_{\rm v}=576$, and the dimension of each visual feature is $d_{\rm v}=1024$. The number of merged features $n_{\rm c}$ varies from $16$ to $64$, and other parameters are consistent with the setup detailed in Section V-A.
Figure 10: The rate-performance curves of different methods on the seven VQA benchmarks relative to the LLaVA-1.5-7B backbone model.
Fig. 10 illustrates the rate-performance curves of different methods on the seven VQA benchmarks relative to the LLaVA-1.5-7B backbone model. The proposed TOFC achieves 70%, 66%, 63%, and 48% reductions in data transmission across different target performance, compared with JPEG, WebP, hyperprior, and ELIC compressions, respectively. The data savings on LLaVA-1.5-7B model are more significant than those on LLaVA-OneVision-7B model, mainly because the input of the former model is only a single image patch. This design reduces the number of elements in the visual features at identical number of merged features per patch $n_{\rm c}$, thus lowering the data transmission overhead of TOFC.
Figure 11: Average inference latency per user request (the bar plot corresponding to the left y-axis) and benchmark score (the line plot corresponding to the right y-axis) under different communication rates on RealWorldQA benchmark with the LLaVA-1.5-7B backbone model.
Fig. 11 shows the average inference latency per user request and the benchmark score under different communication rates on RealWorldQA benchmark with the LLaVA-1.5-7B backbone model. The device-only inference is not incorporated as the LLaVA-1.5 series does not include a model smaller than 7B that is suitable for on-device deployment. The proposed TOFC achieves 47%, 47%, 37%, and 58% reductions in system latency at low communication rate and similar benchmark scores, compared with JPEG, WebP, hyperprior, and ELIC compressions, respectively. Compared to server-only inference, TOFC reduces the end-to-end latency by 68% to 85% across various communication rates while maintaining a 96.5% benchmark score. The reduction in server-side computation time is less prominent for LLaVA-1.5-7B model, whose visual tokens constitute a smaller percentage of the total input tokens, compared with LLaVA-OneVision-7B model. However, the decreased data size of visual features minimizes the data transmission time, and the system latency reduction is thus still significant.
@@ V-C1 Feature Merging Methods
Methods # tokens AI2D MMBench MME MMStar RealWorldQA Avg.
Backbone model [7] 729 82.58 82.13 1584.6 61.80 68.89 100.00%
Proposed FMM 16 74.61 78.78 1540.8 50.33 63.92 91.55%
FMM w/o merging 16 73.83 77.06 1484.4 48.27 61.18 88.77%
VisionZip [28] 16 70.85 74.23 1484.8 46.86 51.37 84.05%
VisionZip merging [28] 16 71.99 75.34 1525.9 49.55 57.25 87.70%
Q-Former [64] 16 68.69 68.04 1379.7 42.83 53.99 80.15%
@@ V-C2 Entropy Coding Methods
To verify the effectiveness of the proposed learnable and selective entropy model with hyperprior in reducing data transmission overhead while maintaining task performance, we conduct an ablation study on alternative entropy coding methods. Specifically, we replace the entropy coding method in the proposed TOFC method with the following three methods. First, we discard the selective entropy model architecture proposed in Section IV-C, retaining only a single entropy model. Second, we substitute the learnable entropy model with a heuristic Laplacian probability model, whose mean and scale are calculated from the statistics of the quantized features in each batch. Linear scaling is applied before quantization, such that the 99.9% percentile of the merged features is mapped to the maximum integer. Third, the learnable factorized density model, originally used for estimating the probability distribution of the hyperprior $\boldsymbol{z}$, is directly employed to model the merged features $\boldsymbol{y}$. For the first and third learnable methods, we train multiple models with varying Lagrange multipliers $\lambda$ and numbers of merged features $n_{\rm c}$ to obtain the rate-performance curve. For the second heuristic method, we adjust the bit length used for quantization and the number of merged features $n_{\rm c}$ to traverse the rate-performance curve.
As shown in Fig. 12, the proposed selective entropy model outperforms alternative entropy coding methods across all data sizes. Compared with a single entropy model, the proposed selective entropy model achieves a 13% to 22% reduction in data transmission while maintaining similar performance. Compared with the heuristic probability model, the proposed entropy model reduces data transmission overhead by approximately 40%, showing that the learnable entropy model provides a more accurate estimation of the statistics of each merged feature. The learnable entropy model without hyperprior delivers poor performance, confirming that the factorized density model alone is insufficient for estimating the probability distribution without the knowledge extracted by the hyperprior $\boldsymbol{z}$, which only constitutes less than 2% of the data transmission and incurs marginal transmission overhead.